Total Production 1999 2000 PriceTotal Production Price Total Production Hotdogs T-Shirts Factories 1000 2000 10 2000 4000 20 100 100 26. The chart above shows total production in an economy and the prices of each good produced in 1999 and 2000. Let 1999 be the base year, and let the market basket for the CPI be I hotdog and I T-shirt. Which of the the following statements is true? a. The inflation rate calculated by using the CPI is lower than the inflation rate calculated by using b. The inflation rate calculated by using the CPI is higher than the inflation rate calculated by using c. The inflation rate calculated by using the CPI is equal to the inflation rate calculated by using the d. the GDP deflator the GDP deflator. GDP deflator. More information is required to compare the inflation rates calculated by using the CPI vs. using the GDP deflator. potlatel 01 -
Answer #1

Answer 26

The correct answer is (b) The Inflation rate calculated by using the CPI is higher than the Inflation rate calculated by using the GDP Deflator

GDP deflator is given by (Nominal GDP/Real GDP)*100

Here Base Year is 1999 and hence Let GDP deflator in 1999 = 100

Lets Calculate Nominal GDP in 2000:

Nominal GDP in 2000 = Sum of Prices in 2000*Quantities of all goods

= 2*2000 + 4*4000 + 100*20 = 22000

Real GDP in 2000 = Sum of Prices (in 1999)*Quantities of all goods (in 2000)

= 1*2000 + 1*4000 + 100*20 = 8000

Hence GDP deflator in 2000 = (22000/8000)*100 = 275

Hence Inflation using GDP deflator = ((GDP deflator in 2000 - GDP deflator in 1999) / (GDP deflator in 1999)) * 100

= ((275-100)/100)*100 = 175

Now Lets Calculate CPI in 2000

CPI in 2000 = (Market value of Basket in 2000/Market value of Basket in 2000) * 100

Basket = 1 Hotdog and 1 Tshirt

So, Market value of Basket in 2000 = 1*2 + 1*4 = 6

Market value of Basket in 1999 = 1*1 + 1*1 = 2

Hence, CPI in 2000 = (6/2)*100 = 300

CPI in 1999(Base Year) = 100

Hence Inflation using CPI = ((CPI in 2000 - CPI in 1999)/(CPI in 1999))*100

= ((300 - 100) / 100) *100 = 200

Hence, Inflation using CPI > Inflation using GDP deflator

Hence Inflation rate calculated by using the CPI is higher than Inflation rate calculated by using the GDP Deflator

Hence the correct answer is (b) The Inflation rate calculated by using the CPI is higher than the Inflation rate calculated by using the GDP Deflator